OpenStudy (anonymous):

Inequality problem: A sign at Pablo's Pizza says that the max number of people it can serve is 225 customers. The restaurant has 3 large tables which seat 12 people at each and x tables which seat 6 people each. What is the inequality that can be used to determine the number of tables for 6 in the restaurant?

OpenStudy (anonymous):

\[3(12)+6x \le 225\] \[36+6x \le 225\] \[6x \le 189\]\[x \le 186/6\] \[x \le 63/2\]

OpenStudy (anonymous):

or \[x \le 31 and 1/2\] (if you divide to get a mixed number)